OpenStudy (brittany16):

Solve the equation. 5t-5/-10=70 i have tried just cant figure it out

OpenStudy (anonymous):

is that (5t-5) / -10 = 70 or 5t - (5 / -10) = 70

OpenStudy (nubeer):

can u put it a little better way with brackets because it's kinda hard to figure the exact question

OpenStudy (brittany16):

no its like 5T-5 over -10=70

OpenStudy (anonymous):

ok, so multiply both sides of the equation by -10 you get 5t-5 = 70 * -10

OpenStudy (nubeer):

-139

OpenStudy (anonymous):

i need sleep sry. 5t-5=-700. good catch ds. night y'all